# Bastani, Bastani, and McLaughlin

### v2

@@ 2.4 Multiarm Characterization
The lack of a well defined treatment effect when comparing multiple treatments complicates the multiarm characterization of $\Pi_{\Lambda}$ ($K>1$) relative to the binary one. The policies in $\Pi_{\Lambda}$ can still be parameterized by the dual variable which uniquely generates them $\xi$ and still form a path of policies from $\pi^{o}$ to the expectation maximizing policy. However, for each unit $n$, the rate of change for each non-zero propensity $\pi^{*}_{n,t}(\xi)$ is impacted by which of the unit’s treatments receive zero propensity. These inactive treatments do not impact the weighted average of treatments which act as a control for each treatment’s relative impact, causing $\pi^{*}_{n,t}(\xi)$ to change in a piece-wise linear manner as $\xi$ increases. To overcome this issue, our general characterization of $\Pi_{\Lambda}$ first determines the value of $\xi$ at which each propensity $\pi^{*}_{n,t}(\xi)$ first hits 0, and then determines the non-zero propensities in terms of the set of zero propensity treatments. Doing so allows us to first identify all of the knot points in the piecewise linear evolution of $\pi^{*}(\xi)$, and then describe $\pi^{*}(\xi)$ in terms of its locally linear form. First, we have: {repeatproposition} Let $(\pi^{\prime},\xi^{\prime},\kappa^{\prime})$ and $(\pi^{\prime\prime},\xi^{\prime\prime},\kappa^{\prime\prime})$ both represent policies in $\Pi_{\Lambda}$ with their accompanying dual variables. Furthermore, for any policy $\pi^{\cdot}\in\Pi_{\Lambda}$ let
@@
$\omega_{n}^{\cdot}=\xi^{\cdot}+N\sum_{t=0}^{K}\frac{\pi_{n,t}^{o}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}\kappa_{n,t}^{\cdot}\left(\mu_{n,t}-\tilde{\mu}_{n}\right).$
@@
If $\omega_{n}^{\prime\prime}\geq\omega_{n}^{\prime}>0$ and $\pi^{\prime}_{n,t}=0$ for some unit $n$ and treatment $t$, then $\pi^{\prime\prime}_{n,t}=0$. Proposition 2.4 ensures that for each $\omega^{*}$ which generates a solution, there exists a unique set $\mathcal{Z}_{n}(\omega^{*})=\{t\mid\pi^{*}_{n,t}(\xi^{*})=0\}$ of treatments which receive no assignments for unit $n$, and that $\mathcal{Z}_{n}(\omega^{\prime})\subseteq\mathcal{Z}_{n}(\omega^{\prime\prime})$ if $\omega^{\prime\prime}\geq\omega^{\prime}>0$ both generate policies in $\Pi_{\Lambda}$. This provides enough structure to uniquely solve for $\kappa^{*}$ in Proposition 2.2 for any $\omega^{*}\in\mathbb{R}_{+}$, generating the full solution to $\Pi_{\Lambda}$ (as $\omega^{*}<0$ induces $\lambda<0$), which can still be uniquely parameterized in terms of $\xi^{*}\geq 0$. {repeattheorem} For each unit $n$, set $\omega_{n}^{0}=0$ and $\hat{\mathcal{Z}}_{n}=\{\}$, and apply this procedure to form $\xi_{n,t}$:
@@
    If $\mu_{n,v}=\mu_{n,u}\ \forall\ v,u\not\in\hat{\mathcal{Z}}_{n}$, set $\xi_{n,t}=\infty\ \forall\ t\not\in\hat{\mathcal{Z}}_{n}$, then all $\xi_{n,t}$ have been formed. Otherwise move to step 2).
@@
    $\xi_{n,t}$ $=\left[1-\sum_{v\not\in\hat{\mathcal{Z}}_{n}}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\left(\mu_{n,v}-\frac{\sum_{j\not\in\hat{\mathcal{Z}}_{n}}\frac{\pi^{o}_{n,j}}{\mu_{n,j}^{2}+\sigma_{n,j}^{2}}\mu_{n,j}}{\sum_{j\not\in\hat{\mathcal{Z}}_{n}}\frac{\pi^{o}_{n,j}}{\mu_{n,j}^{2}+\sigma_{n,j}^{2}}}\right)^{2}\right]\frac{\omega_{n}^{0}}{Q_{n}}$
@@
Let $\mathcal{Z}_{n}(\xi)=\{t|\xi_{n,t}\leq\xi\}.$
For all $K\geq 1$, we have $\Pi_{\Lambda}=\bigcup_{\xi\geq 0}\left\{\pi^{*}(\xi)\right\},$ where $\pi^{*}_{n,t}(\xi)=0\ \forall\ t\in\mathcal{Z}_{n}(\xi)$ while
$\pi^{*}_{n,t}(\xi)=\pi^{o}_{n,t}+\left(\frac{\frac{\pi_{n,t}^{o}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}}{\sum_{v\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi_{n,v}^{o}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}}\right)\left(\sum_{v\in\mathcal{Z}_{n}(\xi)}\pi^{o}_{n,v}\right)+\left(\mu_{n,t}-\frac{\sum_{v\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\mu_{n,v}}{\sum_{v\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}}\right)\frac{N\omega_{n}^{*}(\xi)\pi^{o}_{n,t}}{2Q_{n}\left(\mu_{n,t}^{2}+\sigma_{n,t}^{2}\right)}$
@@
$\omega_{n}^{*}(\xi)$ $=\frac{\xi-\frac{2}{N}\left(\sum_{t\in\mathcal{Z}_{n}(\xi)}\pi_{n,t}^{o}\left(\mu_{n,t}-\frac{\sum_{v\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\mu_{n,v}}{\sum_{v\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}}\right)\right)}{1-\sum_{t\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi^{o}_{n,t}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}\left(\mu_{n,t}-\frac{\sum_{v\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\mu_{n,v}}{\sum_{v\not\in\mathcal{Z}_{n}(\xi)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}}\right)^{2}}Q_{n}$
for all $t\not\in\mathcal{Z}_{n}(\xi)$.
@@
To compute $\mathcal{Z}_{n}(\xi)$ for each unit, we iteratively compute which treatment’s propensity will first hit zero for each unit, remove that treatment and continue. An added benefit to the computation and analysis of $\pi^{*}(\xi)$ is that each unit’s treatment propensities can be calculated independently. We see that as $\xi$ increases, $\pi^{*}(\xi)$ moves from the observational policy to the expectation maximizing policy by iteratively eliminating suboptimal treatment arms for each unit.
@@
The non-zero propensities of $\pi^{*}(\xi)$ are determined by a term which redistributes zero propensity treatments and a linear term similar to the binary characterization. This linear term replaces the CATE with a multi-treatment version which compares the expected outcome of unit $n$ in response to treatment $t$ to a weighted average of the expected outcome of the other active treatments. The variance of the conditional treatment effect is then replaced by $Q_{n}^{-1}\frac{\pi^{o}_{n,t}}{\mu_{n,t}^{2}+\sigma^{2}_{n,t}}$. The treatment specific term here, which also determines how zero propensity treatments are redistributed and the weighted mean each treatment is measured against, is once again related to IPW’s evaluation:
$\frac{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}{\pi^{o}_{n,t}}=\mathbb{E}\left[\left(\mathbf{1}\left[T_{n}^{o}=t\right]\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right)^{2}\right]=\mathrm{Var}\left(\mathbf{1}\left[T_{n}^{o}=t\right]\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right)+\mathbb{E}\left[\mathbf{1}\left[T_{n}^{o}=t\right]\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right]^{2}.$
@@
Thus, as $\xi$ increases, $\pi^{*}(\xi)$ evolves similarly to the binary treatment case; it computes a conditional average treatment effect for each treatment against the other active treatments, and then changes each unit’s policy to increase the policies impact while introducing minimal variance. Along the way, it redistributes inactive treatments in a way that maintains the continuity of $\pi^{*}(\xi)$.
2.5 Connection to Pareto Frontier
@@
Extracting $\Pi_{\mathrm{Opt}}$ from $\Pi_{\Lambda}$ involves comparing the expected improvement and expected $z$-score achieved by policies in $\Pi_{\Lambda}$. Luckily, the policies in $\Pi_{\Lambda}$ can be continuously parameterized via the dual variable $\xi$, which relates to both the evolution of the expected improvement and expected $z$-score of $\pi^{*}(\xi)$. At a high level, the expected improvement is strictly increasing as $\pi^{*}(\xi)$ moves from the observational policy toward the expectation maximizing policy, while the expected $z$-score is constant for small $\xi$, and then strictly decreasing after any unit has a treatment reach zero propensity. Thus, we can characterize $\Pi_{\mathrm{Opt}}$ as the region of $\Pi_{\Lambda}$ in which expected improvements are increasing while expected $z$-score is decreasing.
First, we show that the expected value of $\pi^{*}(\xi)$ increases with $\xi$ as the policy moves towards the expectation maximizing policy. Let $\xi_{\max}=\underset{\xi\geq 0}{\mathrm{argmin}}\{\xi\mid\mathcal{Z}_{n}(\xi)=\lim_{\xi\to\infty}\mathcal{Z}_{n}(\xi)\ \forall\ n\in[N]\}$ be the largest $\xi$ for which a treatment propensity first reaches 0. For $\xi\geq\xi_{\max}$ we know $\pi^{*}(\xi)$ recovers the expectation maximizing policy. Up until that point, $\pi^{*}(\xi)$ increases the variance of its policy evaluation to improve its expectation as $\xi$ increases. {repeatproposition} $\tau(\pi^{*}(\xi))$ is increasing in $\xi$ for $\xi\in[0,\xi_{max}$). Next, the expected $z$-score of $\pi^{*}(\xi)$ (weakly) decreases as $\xi$ increases, meaning $\Pi_{\Lambda}$ maps a trade-off between expected policy improvement and the expected $z$-score a policy achieves when evaluated by IPW. Let $\xi_{\min}=\underset{\xi\geq 0}{\mathrm{argmin}}\{\xi\mid\min_{n\in[N],t\in\{1,\ldots,K\}}\pi_{n,t}^{*}(\xi)=0\}$ be the smallest $\xi$ for which some unit is never assigned one of the treatments. For $\xi\in[0,\xi_{\min})$, $\pi^{*}(\xi)$ produces strictly positive propensities and maintains a constant expected $z$-score under evaluation by IPW. Beyond $\xi_{\min}$, increasing the expected improvement of $\pi^{*}(\xi)$ requires a decrease in the expected $z$-score of $\pi^{*}(\xi)$ as it can no longer decrease (low mean) treatment propensities which have already been set to zero. {repeatproposition} $\mathbb{E}[Z(\pi^{*}(\xi))]$ is constant in $\xi$ for $\xi\in(0,\xi_{min})$ and strictly decreasing in $\xi$ for $\xi\in(\xi_{\min},\xi_{\max})$.
As we already know, $\Pi_{\text{Opt}}\subseteq\Pi_{\Lambda}$ by Proposition 2.2; thus, Propositions $2.5$ allow us to characterize $\Pi_{\text{Opt}}$ as the policies in $\Pi_{\Lambda}$ which strictly trade-off expected policy improvement and expected $z$-score when evaluated by IPW. {repeattheorem} $\Pi_{\mathrm{Opt}}=\left\{\pi^{*}(\xi)|\xi\in[\xi_{\min},\xi_{\max}]\right\}$. Theorem 2.5 shows that there is some small amount of “free lunch” in policy optimization. At least one treatment from a single unit can be eliminated with some amount of internal propensity redistribution to form the policy $\pi^{*}(\xi_{\min})$, which achieves positive expected improvement and maximal expected $z$-score. Proposing policies which eliminate more treatments to achieve a higher expected improvement necessarily requires accepting policy changes which deteriorate the significance of the improvement when evaluated on a held out sample. Therefore, by Theorem 2.4, practitioners with knowledge of the counterfactual outcomes distribution can select a policy that balances their preferences for expected improvement and significance when evaluated out of sample.

### v3

@@ 4.1 Superset of the Pareto Frontier
The lack of a well defined treatment effect when comparing multiple treatments complicates the multiarm characterization of $\Pi_{\Lambda}$ ($K>1$) relative to the single-treatment setting. The policies in $\Pi_{\Lambda}$ can still be parameterized by the dual variable which uniquely generates them $\zeta$ and still form a path of policies from $\pi^{o}$ to the expectation maximizing policy. However, for each unit $n$, the rate of change for each non-zero propensity $\pi^{*}_{n,t}(\zeta)$ is impacted by which of the unit’s treatments receive zero propensity. These inactive treatments do not impact the weighted average of treatments which act as a control for each treatment’s relative impact, causing $\pi^{*}_{n,t}(\zeta)$ to change in a piece-wise linear manner as $\zeta$ increases. To overcome this issue, our general characterization of $\Pi_{\Lambda}$ first determines the value of $\zeta$ at which each propensity $\pi^{*}_{n,t}(\zeta)$ first hits 0, and then determines the non-zero propensities in terms of the set of zero propensity treatments. Doing so allows us to first identify all of the knot points in the piecewise linear evolution of $\pi^{*}(\zeta)$, and then describe $\pi^{*}(\zeta)$ in terms of its locally linear form. First, we have: {repeatproposition} Let $(\pi^{\prime},\zeta^{\prime},\kappa^{\prime})$ and $(\pi^{\prime\prime},\zeta^{\prime\prime},\kappa^{\prime\prime})$ both represent policies in $\Pi_{\Lambda}$ with their accompanying dual variables. Furthermore, for any policy $\pi^{\cdot}\in\Pi_{\Lambda}$ let
@@
$\omega_{n}^{\cdot}=\zeta^{\cdot}+N\sum_{t=0}^{K}\frac{\pi_{n,t}^{o}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}\kappa_{n,t}^{\cdot}\left(\mu_{n,t}-\tilde{\mu}_{n}\right).$
@@
If $\omega_{n}^{\prime\prime}\geq\omega_{n}^{\prime}>0$ and $\pi^{\prime}_{n,t}=0$ for some unit $n$ and treatment $t$, then $\pi^{\prime\prime}_{n,t}=0$. Proposition 4.1 ensures that for each $\omega^{*}$ which generates a solution, there exists a unique set $\mathcal{Z}_{n}(\omega^{*})=\{t\mid\pi^{*}_{n,t}(\zeta^{*})=0\}$ of treatments which receive no assignments for unit $n$, and that $\mathcal{Z}_{n}(\omega^{\prime})\subseteq\mathcal{Z}_{n}(\omega^{\prime\prime})$ if $\omega^{\prime\prime}\geq\omega^{\prime}>0$ both generate policies in $\Pi_{\Lambda}$. This provides enough structure to uniquely solve for $\kappa^{*}$ in Proposition 2.2 for any $\omega^{*}\in\mathbb{R}_{+}$, generating the full solution to $\Pi_{\Lambda}$ (as $\omega^{*}<0$ induces $\lambda<0$), which can still be uniquely parameterized in terms of $\zeta^{*}\geq 0$. {repeattheorem} For each unit $n$, set $\omega_{n}^{0}=0$ and $\hat{\mathcal{Z}}_{n}=\{\}$, and apply this procedure to form $\zeta_{n,t}$:
@@
    If $\mu_{n,v}=\mu_{n,u}\ \forall\ v,u\not\in\hat{\mathcal{Z}}_{n}$, set $\zeta_{n,t}=\infty\ \forall\ t\not\in\hat{\mathcal{Z}}_{n}$, then all $\zeta_{n,t}$ have been formed. Otherwise move to step 2).
@@
    $\zeta_{n,t}$ $=\left[1-\sum_{v\not\in\hat{\mathcal{Z}}_{n}}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\left(\mu_{n,v}-\frac{\sum_{j\not\in\hat{\mathcal{Z}}_{n}}\frac{\pi^{o}_{n,j}}{\mu_{n,j}^{2}+\sigma_{n,j}^{2}}\mu_{n,j}}{\sum_{j\not\in\hat{\mathcal{Z}}_{n}}\frac{\pi^{o}_{n,j}}{\mu_{n,j}^{2}+\sigma_{n,j}^{2}}}\right)^{2}\right]\frac{\omega_{n}^{0}}{Q_{n}}$
@@
Let $\mathcal{Z}_{n}(\zeta)=\{t|\zeta_{n,t}\leq\zeta\}.$
For all $K\geq 1$, we have $\Pi_{\Lambda}=\bigcup_{\zeta\geq 0}\left\{\pi^{*}(\zeta)\right\},$ where $\pi^{*}_{n,t}(\zeta)=0\ \forall\ t\in\mathcal{Z}_{n}(\zeta)$ while
$\pi^{*}_{n,t}(\zeta)=$ $\pi^{o}_{n,t}+\left(\frac{\frac{\pi_{n,t}^{o}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}}{\sum_{v\not\in\mathcal{Z}_{n}(\zeta)}\frac{\pi_{n,v}^{o}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}}\right)\left(\sum_{v\in\mathcal{Z}_{n}(\zeta)}\pi^{o}_{n,v}\right)$
$+\left(\mu_{n,t}-\frac{\sum_{v\not\in\mathcal{Z}_{n}(\zeta)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\mu_{n,v}}{\sum_{v\not\in\mathcal{Z}_{n}(\zeta)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}}\right)\frac{N\omega_{n}^{*}(\zeta)\pi^{o}_{n,t}}{2Q_{n}\left(\mu_{n,t}^{2}+\sigma_{n,t}^{2}\right)}$
@@
$\omega_{n}^{*}(\zeta)$ $=\frac{\zeta-\frac{2}{N}\left(\sum_{t\in\mathcal{Z}_{n}(\zeta)}\pi_{n,t}^{o}\left(\mu_{n,t}-\frac{\sum_{v\not\in\mathcal{Z}_{n}(\zeta)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\mu_{n,v}}{\sum_{v\not\in\mathcal{Z}_{n}(\zeta)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}}\right)\right)}{1-\sum_{t\not\in\mathcal{Z}_{n}(\zeta)}\frac{\pi^{o}_{n,t}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}\left(\mu_{n,t}-\frac{\sum_{v\not\in\mathcal{Z}_{n}(\zeta)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}\mu_{n,v}}{\sum_{v\not\in\mathcal{Z}_{n}(\zeta)}\frac{\pi^{o}_{n,v}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}}\right)^{2}}Q_{n}$
@@
for all $t\not\in\mathcal{Z}_{n}(\zeta)$. To compute $\mathcal{Z}_{n}(\zeta)$ for each unit, we iteratively compute which treatment’s propensity will first hit zero for each unit, remove that treatment and continue. An added benefit to the computation and analysis of $\pi^{*}(\zeta)$ is that each unit’s treatment propensities can be calculated independently. We see that as $\zeta$ increases, $\pi^{*}(\zeta)$ moves from the observational policy to the expectation maximizing policy by iteratively eliminating suboptimal treatment arms for each unit.
@@
The non-zero propensities of $\pi^{*}(\zeta)$ are determined by a term which redistributes zero propensity treatments and a linear term similar to the single-treatment characterization. This term replaces $\tau_{n}$ with a multi-treatment version which compares the expected outcome of unit $n$ in response to treatment $t$ to a weighted average of the expected outcome of the other active treatments. The variance of the conditional treatment effect is then replaced by $Q_{n}^{-1}\frac{\pi^{o}_{n,t}}{\mu_{n,t}^{2}+\sigma^{2}_{n,t}}$. The treatment specific term here, which also determines how zero propensity treatments are redistributed as well as the weighted mean each treatment is measured against, is once again related to IPW’s evaluation:
$\frac{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}{\pi^{o}_{n,t}}=\mathbb{E}\left[\left(\mathbbm{1}\left[T_{n}^{o}=t\right]\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right)^{2}\right]=\mathrm{Var}\left(\mathbbm{1}\left[T_{n}^{o}=t\right]\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right)+\mathbb{E}\left[\mathbbm{1}\left[T_{n}^{o}=t\right]\frac{Y_{n}^{o}}{\pi^{o}_{n,T_{n}^{o}}}\right]^{2}.$
@@
Thus, as $\zeta$ increases, $\pi^{*}(\zeta)$ evolves similarly to the single-treatment case; it compares each treatment against the other active treatments, and then changes each unit’s policy to increase the policies impact while introducing minimal variance to an estimate of the policy’s improvement.
4.2 Structure of the Pareto Frontier
@@
As in the single-treatment setting, the expected improvement is strictly increasing as $\pi^{*}(\zeta)$ moves from the observational policy toward the expectation-maximizing policy, while the expected $z$-score is constant for small $\zeta$, and then strictly decreasing after any single unit has a treatment reach zero propensity. Define $\tau(\zeta)=\tau(\pi^{*}(\zeta))$, $s^{2}(\zeta)=s^{2}(\pi^{*}(\zeta))$, and $Z(\zeta)=\mathbb{E}[Z(\pi^{*}(\zeta))]=\frac{\tau(\zeta)}{s^{2}(\zeta)}$; also,
$\zeta_{\text{max}}$ $=\operatorname*{\arg\min}_{\zeta\geq 0}\{\zeta\mid\mathcal{Z}_{n}(\zeta)=\lim_{\zeta\to\infty}\mathcal{Z}_{n}(\zeta)\ \forall\ n\in[N]\}$
$\zeta_{\text{min}}$ $=\operatorname*{\arg\min}_{\zeta\geq 0}\{\zeta\mid\min_{n\in[N],t\in\{1,\ldots,K\}}\pi_{n,t}^{*}(\zeta)=0\}.$
In other words, $\zeta_{\text{max}}$ is the largest $\zeta$ for which a treatment propensity first reaches 0. For $\zeta\geq\zeta_{\text{max}}$ we know $\pi^{*}(\zeta)$ recovers the expectation maximizing policy; up until that point, $\pi^{*}(\zeta)$ increases the variance of its policy evaluation to improve its expectation as $\zeta$ increases. Similarly, $\zeta_{\text{min}}$ is the smallest $\zeta$ for which some unit is never assigned one of the treatments. {repeatproposition} $\tau(\zeta)$ is strictly increasing for $\zeta\in(0,\zeta_{\text{max}})$. {repeatproposition} $Z(\zeta)$ is constant in $\zeta$ for $\zeta\in(0,\zeta_{\text{min}})$ and strictly decreasing for $\zeta\in(\zeta_{\text{min}},\zeta_{\text{max}})$. {repeattheorem} The Pareto frontier is $\Pi_{\mathrm{Opt}}=\{\pi^{*}(\zeta)\mid\zeta\in[\zeta_{\text{min}},\zeta_{\text{max}}]\}$. This result mirrors Corollary 1 in the single-treatment case. The main difference is that only a single unit-treatment pair needs to bind to zero at $\zeta_{\text{min}}$; thus, all units may have stochastic treatments.
4.3 Structure of Policies on the Pareto Frontier
Next, we analyze how policies on the Pareto frontier assign treatments that trade off statistically significance vs. expected improvement. At a high level, as $\zeta$ increases, for each unit, policies along the frontier increase the propensity assigned to treatments with expected value above some unit-specific threshold, while decreasing the propensity assigned to the remaining treatments. In addition, a treatment’s propensity will remain closer to that of the observational policy if that treatment’s counterfactual outcome is often far from 0. Thus policies on the frontier react conservatively to variance, choosing to adjust the propensities of treatments leading to certain outcomes sooner than uncertain ones.
Our first result shows that policies along the frontier decrease treatment propensities below the observational policy in order of their expected impact. {repeatproposition} For unit $n$, if $t$ and $t^{\prime}$ are treatments with $\mu_{n,t}>\mu_{n,t^{\prime}}$ and $\frac{\pi_{n,t^{\prime}}^{*}(\zeta)}{\pi_{n,t^{\prime}}^{o}}>1$, then
$\frac{\pi_{n,t}^{*}(\zeta)}{\pi_{n,t}^{o}}>1.$
This result follows a simple logic: for any policy that does not obey this ordering, we can select a pair of offending treatments and contract their propensities towards the observational policy to both increase expected performance and decrease the standard error of the IPW estimator. As such we can understand policies on the frontier as setting a reservation value for each unit; increasing the frequency of treatments in the subgroup that outperform this reservation value while decreasing the frequency of treatments which underperform the reservation value.
However, we also show that policies along the frontier adjust treatments whose counterfactual outcomes are far from 0 more conservatively, meaning a treatment whose average impact is worse than an alternate may be assigned relatively more. {repeatproposition} For unit $n$, if $t$ and $t^{\prime}$ are treatments with $\mu_{n,t}>\mu_{n,t^{\prime}}$ and $\frac{\pi^{*}_{n,t^{\prime}}(\zeta)}{\pi^{o}_{n,t^{\prime}}}>\frac{\pi^{*}_{n,t}(\zeta)}{\pi^{o}_{n,t}}$, then
  • (1)
    $\frac{\pi^{*}_{n,t^{\prime}}(\zeta)}{\pi^{o}_{n,t^{\prime}}}>1\implies\underbrace{\mathbb{E}[Y_{n,t}^{2}]}_{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}>\underbrace{\mathbb{E}[Y_{n,t^{\prime}}^{2}]}_{\mu_{n,t^{\prime}}^{2}+\sigma_{n,t^{\prime}}^{2}}$
  • (2)
    $\frac{\pi^{*}_{n,t^{\prime}}(\zeta)}{\pi^{o}_{n,t^{\prime}}}<1\implies\underbrace{\mathbb{E}[Y_{n,t}^{2}]}_{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}<\underbrace{\mathbb{E}[Y_{n,t^{\prime}}^{2}]}_{\mu_{n,t^{\prime}}^{2}+\sigma_{n,t^{\prime}}^{2}}$
Note that under Proposition 4.3, the conditions of Proposition 4.3 can only occur if the frontier policy increases both treatments’ propensities above that of the observational policy, or decreases both below that of the observational policy. Thus condition (1) says that if the frontier policy increases both treatments, it may increase the lower mean treatment relatively more if its average outcome is closer to zero, while condition (2) says that if the frontier policy decreases both treatments, it may decrease the lower mean treatment relatively less if its average outcome is closer to zero. Re-ranking treatments in this way reduces the variance of the IPW estimator as the estimator scales observed outcomes in the test set with a magnitude that increases with the difference between the frontier policy and the observational policy. If outcomes are far from zero, this scaling can introduce excessive variance that destroys the potential of the IPW estimator to achieve significance, leading policies on the Pareto frontier to adjust treatments with counterfactual outcomes close to zero more than other treatments.
4.4 Statistical Power
Finally, we compute the maximum expected z-score any policy can achieve, which is the analog of Corollary 3.3 in the general setting. {repeatproposition} Let
$\pi_{n,t}^{\dagger}$ $=\frac{\mathbb{E}\left[\mathbbm{1}[T_{n}^{o}=t]\mathbb{E}[Y_{nT_{n}^{o}}^{2}\mid T_{n}^{o}]^{-1}\right]}{\mathbb{E}\left[\mathbb{E}\left[Y_{nT_{n}^{o}}^{2}\mid T_{n}^{o}\right]^{-1}\right]}=\frac{\frac{\pi_{n,t}^{o}}{\mu_{n,t}^{2}+\sigma_{n,t}^{2}}}{\sum_{v=0}^{K}\frac{\pi_{n,v}^{o}}{\mu_{n,v}^{2}+\sigma_{n,v}^{2}}},$ $T_{n}^{\dagger}$ $\sim\mathrm{Categorical}(\pi_{n}^{\dagger}),$
then
$\max_{\pi\in(\Delta^{K})^{N}}\mathbb{E}[Z(\pi)]=\sqrt{\sum_{n=1}^{N}\frac{\mathrm{Var}(\mu_{n,T^{\dagger}_{n}})}{\mathbb{E}[\sigma^{2}_{n,T^{\dagger}_{n}}]+\mathbb{E}[\mu_{n,T^{\dagger}_{n}}]^{2}}}.$
Intuitively, each unit’s contribution increases with the treatment effect for that unit and decreases with the variance in its treatment effect and with the magnitude of the mean outcomes.